What mass of calcium oxide, CaO, forms when 28g calcium carbonate, CaCO3, decomposes on heating?
1
Expert's answer
2020-09-24T11:20:34-0400

CaCO3 = CaO+CO2

n (CaO)=n(CaCO3)

n=m/M

M(CaCO3)=100 g/mol

M(CaO)=56 g/mol

n(CaCO3)=28/100=0.28 mol

n (CaO)=n(CaCO3)=0.28 mol

m (CaO)=n x M = 0.28 x 56 = 15.7 g